# ActSonic: Recognizing Everyday Activities from Inaudible Acoustic Waves Around the Body

### v2

@@ 4. Deep Learning Framework
To estimate human activities from acoustic sensing data in ActSonic, we design a self-supervised deep learning framework which is illustrated in Fig. 4. This framework leverages the unlabelled acoustic data to create a pre-trained encoder. Then, we use this encoder to create an activity recognition pipeline.
@@ 4.1.1. Pretraining Task
The pre-training approach to learning representation from the unlabelled data is to perturb the sliding window of acoustic flow (described in 3.2) with binary mask and reconstruct the original window using the autoencoder depicted in the pre-training task segment of Fig. 4. Here, the binary mask is constructed in a way such that $m\%$ of each channel of the sliding window is set to 00. In the case of ActSonic, the numerical value $m$ of this mask percentage is randomly chosen from the range $15-20\%$, and the number of patches is randomly chosen from the range $1$ to $4$. As illustrated in Fig. 4, the masked window goes through a ResNet18 (He et al., 2015) encoder. The embedded representation from the encoder is then fed through a decoder network, essentially a transpose convolution network. This customized transpose convolution, or deconvolution, network takes feature maps from the ResNet18 as input and generates a three-dimensional matrix of the shape of the input sliding window. We calculate the Mean Squared Error (MSE) between each pixel in the reconstructed and original sliding window as the loss function for this autoencoder architecture.
@@ 4.1.2. Fine-tuning
The aforementioned ResNet18 encoder of the pre-training pipeline (detailed in Subsec. 4.1.1) learns the representation of active acoustic data via self-supervision. This ResNet18 encoder with learned weights serves as the feature extraction pipeline in the fine-tuning phase. We design an activity recognition architecture (depicted in the fine-tuning segment of Fig. 4) comprising of pre-trained ResNet18 encoder followed by a fully connected classifier layer. We apply average pooling on the spatial axis of the feature map extracted by the encoder and feed it to the fully connected layer. The fully connected classifier network is a feedforward neural network with batch normalization (Ioffe and Szegedy, 2015), Leaky ReLU activation (Xu et al., 2015), and dropout (Srivastava et al., 2014) in between. We set the number of neurons in the last layer equal to the number of activity classes and perform a softmax operation to output a probability distribution.
@@ 5.2. Participants and Study Schedule
Based on insights from a pilot feasibility study, human activity patterns vary throughout the day. For instance, activities such as tooth brushing are more likely to occur in the morning or after dinner. Accordingly, we scheduled user studies to capture activity data across different parts of participants’ daily routines. As shown in Fig. 5, each point represents the start time of a study session.
@@ 5.5. Peer-reviewed Data Annotation Protocol
To provide annotations for active acoustic data via ground-truth egocentric video data, we utilized the ANU-CVML Video Annotation Tool (Vidat) (Zhang et al., 2020) to annotate all ground-truth egocentric video data with action annotations. Subsequently, we synchronized the timestamps of acoustic and video data using a clapping action with a distinct acoustic signature. We then developed separate postprocessing scripts to align video annotations with acoustic echo profiles. To ensure accurate annotation of activities in a naturalistic setting, we implemented a peer-reviewed annotation process. In this procedure, one annotator from the research team labeled the data, while another researcher independently reviewed the annotations and provided feedback. After a phase of revision and approval from the reviewer, the annotated labels were incorporated into the ActSonic dataset.
@@ 6.2. Evaluation of User Independent Model on Unconstrained Sessions
6.3. Power Signature of the System
Our ActSonic system initially consumed $577.8$ mW with the first prototype employing Teensy 4.1. Substituting this microcontroller with a low-power nRF52840 significantly reduced power consumption. We integrated the original speakers and microphones from the initial prototype into the second one, adjusting the gain to ensure identical sound pressure levels (SPL) for both, maintaining emission consistency. The power signature, measured using a CurrentRanger (Chucri A. Kardous and Peter B. Shaw, [n. d.]), displayed an average operation of $96.5$ mW (4.04.04.04.02 V, $24.0$ mA) while saving all data to the SD card. Furthermore, we conducted long-term stability testing, and the prototype operated continuously for $11.3$ hours using a $290$ mAh $3.7$ V Li-Po battery. This configuration enables a full-day operation on commodity smart glasses or AR glasses. For instance, Google Glass, equipped with a $570$ mAh battery, can support the ActSonic sensing system for over $21$ hours if the activity recognition pipeline is the only active process.
6.4. Latency and System Overhead of ActSonic on Mobile Platform
@@
Table 1. ActSonic ResNet18 model latency and system overhead on Google Pixel 7 android mobile platform
We evaluated the latency and overhead of the ActSonic system on the Google Pixel 7 mobile platform. Table 1 presents the inference time and various parameters of the ActSonic ResNet18 model evaluated on an unconstrained dataset with the same protocol described in 6.2. Initially, we generated lightweight mobile models from both the original and the 8-bit integer quantized versions of the ResNet18 model using PyTorch Mobile. Subsequently, we conducted inference time benchmarks using a two-second sliding window on a Google Pixel 7 Android phone, performing inference on 1000 samples. The mean inference time is detailed in Table 1. The ActSonic performance on mobile devices was evaluated by transmitting the acoustic data to a Pixel 7 for inference. The BLE status was set to ”connected” since it continuously streams data to mobile devices.
@@ 8.1. Visual Explanation of Learned Parameters of ActSonic Model through Saliency Analysis
To address this, we conducted saliency analysis using Gradient-weighted Class Activation Mapping (Grad-CAM) (Selvaraju et al., 2019) to visualize the ResNet18 encoder’s final convolutional layer. Overlaying feature heatmaps on acoustic flow or differential echo profile channels confirm the model’s ability to capture the movements of body parts and areas related to each activity. Figure 11 demonstrates Grad-CAM (Selvaraju et al., 2019) for all activities in the dataset.
@@ 8.3. Model Quantization and MCU Inference
To explore the feasibility of deploying our framework on glasses, we implemented the model pipeline on the MAX78002 microcontroller unit (MCU), leveraging its built-in ultra-low-power CNN accelerator. Initially, we quantized the model by converting high-precision floating-point model parameters to 8-bit integers, a necessary step for deployment on the MAX78002 MCU. Subsequently, we generated a C program for the quantized model inference using the ai8x (AI, [n. d.]) library provided by the MCU manufacturer. Due to hardware constraints, certain adjustments were made to the model pipeline for compatibility. Notably, 2D convolution kernel sizes were limited to $(1\times 1)$ or $(3\times 3)$, with fixed stride size at $(1\times 1)$. Additionally, the fully connected layer was capped at a maximum of $1024$ input neurons on the chip. Although we successfully ran the ActSonic model on the MAX78002 MCU, the computation for echo profile calculation resulted in a slower frame rate than anticipated for real-time inference.
@@ 8.4. Robustness to Ambient Noise
The ActSonic system utilizes active acoustic sensing to detect daily activities by monitoring body motions. We evaluated its resilience to environmental noise across $19$ participants’ homes during the studies. Despite varying environmental factors like HVAC, running water, TV, and ambient sounds, detailed in Sec. 6, ActSonic maintains consistent performance independent of environment settings. This resilience stems from its reliance on ultrasonic frequencies ($18$ KHz to $24.5$ KHz), which surpass most environmental noise sources (recorded at frequencies below $7.5$ KHz). For instance, noises like cafe chatter ($63.8$ dB), roadside traffic ($69.0$ dB), and loud music ($71.5$ dB) fell below ActSonic’s operational range. Additionally, any overlapping frequencies in this range would result in significantly stronger signals near the eyeglasses, reinforcing ActSonic’s robustness to environmental acoustic noise.
@@ 8.5. Privacy Preservation
8.6. Health Implication
ActSonic emits FMCW-encoded ultrasonic waves for active acoustic sensing. To assess health implications, we measured the transmitted signal intensity using a CDC-provided mobile app (LowPowerLab, [n. d.]). The resultant intensity is $68$ dB(A), well below the $85$ dB limit set by NIOSH (Murphy and Franks, 2002). Research (Moyano et al., 2022) on MHz range ultrasonic exposure suggests muscle tissue discomfort. However, ActSonic operates in the KHz range just above the audible threshold, with no reported issues in this range. Future investigations will explore potential audibility among animals and children despite its inaudibility to adults for long-term usage.
@@ 8.7. Potential Real-world Application
8.7.1. Behavioral Journaling
We can utilize this system to journal various everyday activities for different purposes. For instance, one crucial step in combating eating disorders is to journal food intake behavior, often recorded manually. Previous eating journaling systems required multiple sensors, had low time resolution (e.g., recognizing a meal every 10 minutes) (Shin et al., 2022), or necessitated training data from a user. In contrast, ActSonic can recognize eating moments at 1 Hz with over 90% F1 score in real-world settings without the need for training data from a new user. This indicates that our system can be immediately deployed to facilitate eating journaling practices.
8.7.2. High-Resolution Behavior Data in the Wild for Health
ActSonic can track 27 everyday activities, many of which are related to health behaviors. Automatically logging these health-related activities can potentially provide researchers and clinical physicians with an opportunity to better understand a user’s activities in the wild for health purposes. For instance, eating and drinking behavior can potentially be used to analyze the user’s eating and drinking routines related to eating disorders and body hydration levels. Brushing teeth, flossing, and rinsing the mouth can be recognized with a very low error rate. This feature can be utilized by dentists to better track patient dental behavior.
@@ 9. Conclusion
This paper introduces ActSonic, a low-power and unobtrusive action recognition system employing acoustic sensing on smart glasses. The extensive experiments involving $19$ participants in real-world settings showcase ActSonic’s adeptness in distinguishing a diverse range of everyday actions across different environments. We envision ActSonic as a straightforward and efficient supplementary modality for egocentric action recognition, addressing concerns regarding privacy.

### v3

@@ 4. DEEP LEARNING FRAMEWORK
To estimate human activities from acoustic sensing data in ActSonic, we design a self-supervised deep learning framework which is illustrated in Figure 4. This framework leverages the unlabelled acoustic data to create a pre-trained encoder. Then, we use this encoder to create an activity recognition pipeline.
@@ 4.1.1. Pretraining Task
The pre-training approach to learning representation from the unlabelled data is to perturb the sliding window of acoustic flow (described in 3.2) with binary mask and reconstruct the original window using the autoencoder depicted in the pre-training task segment of Figure 4. Here, the binary mask is constructed in a way such that $m\%$ of each channel of the sliding window is set to 00. In the case of ActSonic, the numerical value $m$ of this mask percentage is randomly chosen from the range $15-20\%$, and the number of patches is randomly chosen from the range $1$ to $4$. As illustrated in Figure 4, the masked window goes through a ResNet18 (He et al., 2015) encoder. The embedded representation from the encoder is then fed through a decoder network, essentially a transpose convolution network. This customized transpose convolution, or deconvolution, network takes feature maps from the ResNet18 as input and generates a three-dimensional matrix of the shape of the input sliding window. We calculate the Mean Squared Error (MSE) between each pixel in the reconstructed and original sliding window as the loss function for this autoencoder architecture.
@@ 4.1.2. Fine-tuning
The aforementioned ResNet18 encoder of the pre-training pipeline (detailed in Subsec. 4.1.1) learns the representation of active acoustic data via self-supervision. This ResNet18 encoder with learned weights serves as the feature extraction pipeline in the fine-tuning phase. We design an activity recognition architecture (depicted in the fine-tuning segment of Figure 4) comprising of pre-trained ResNet18 encoder followed by a fully connected classifier layer. We apply average pooling on the spatial axis of the feature map extracted by the encoder and feed it to the fully connected layer. The fully connected classifier network is a feedforward neural network with batch normalization (Ioffe and Szegedy, 2015), Leaky ReLU activation (Xu et al., 2015), and dropout (Srivastava et al., 2014) in between. We set the number of neurons in the last layer equal to the number of activity classes and perform a softmax operation to output a probability distribution.
@@ 5.2. Participants and Study Schedule
Based on insights from a pilot feasibility study, human activity patterns vary throughout the day. For instance, activities such as tooth brushing are more likely to occur in the morning or after dinner. Accordingly, we scheduled user studies to capture activity data across different parts of participants’ daily routines. Users provided their preferred study date and time slots for each part of the day (morning, afternoon, and evening as defined in Figure 5). As shown in Figure 5, each point represents the start time of a study session, and the respective users were randomly assigned based on their preferences.
@@ 5.5. Peer-reviewed Data Annotation Protocol
To provide annotations for active acoustic data via ground-truth egocentric video data, we utilized the ANU-CVML Video Annotation Tool (Vidat) (Zhang et al., 2020) to annotate all ground-truth egocentric video data with action annotations. Subsequently, we synchronized the timestamps of the acoustic and video data using a clapping action at the beginning of the recording, which had a distinct acoustic signature. We then developed separate postprocessing scripts to align video annotations with acoustic echo profiles. To ensure accurate annotation of activities in a naturalistic setting, we implemented a peer-reviewed annotation process. In this procedure, one annotator from the research team labeled the data, while another researcher independently reviewed the annotations and provided feedback. After a phase of revision and approval from the reviewer, the annotated labels were incorporated into the ActSonic dataset.
@@ 6.2. Evaluation of User Independent Model on Unconstrained Sessions
6.3. Power Signature of the ActSonic Sensing System on Eyeglasses Form Factor
ActSonic system initially consumed $577.8$ mW for collecting data with the first prototype employing Teensy 4.1. Substituting this microcontroller with a low-power nRF52840 significantly reduced power consumption. We integrated the original speakers and microphones from the initial prototype into the second one, adjusting the gain to ensure identical sound pressure levels (SPL) for both, maintaining emission consistency. The power signature, measured using a CurrentRanger (LowPowerLab, [n. d.]), displayed an average operation of $96.5$ mW (4.04.04.04.02 V, $24.0$ mA) while saving all data to the SD card. Furthermore, we conducted long-term stability testing, and the prototype operated continuously for $11.3$ hours using a $290$ mAh $3.7$ V Li-Po battery. This configuration enables a full-day operation on commodity smart glasses or AR glasses. For instance, Google Glass, equipped with a $570$ mAh battery, can support the ActSonic sensing system for over $21$ hours if the activity recognition pipeline is the only active process.
6.4. Latency and System Overhead of ActSonic Inference Pipeline on Mobile Platform
@@
Table 1. ActSonic ResNet18 model latency and system overhead on Google Pixel 7 android mobile platform.
We evaluated the latency and overhead of the ActSonic inference pipeline on the Google Pixel 7 mobile platform. Table 1 presents the inference time and various parameters of the ActSonic ResNet18 model evaluated on an unconstrained dataset with the same protocol described in 6.2. Initially, we generated lightweight mobile models from both the original and the 8-bit integer quantized versions of the ResNet18 model using PyTorch Mobile. Subsequently, we conducted inference time benchmarks using a two-second sliding window on a Google Pixel 7 Android phone, performing inference on 1000 samples. The mean inference time is detailed in Table 1. The ActSonic performance on mobile devices was evaluated by transmitting the acoustic data to a Pixel 7 for inference. The BLE status was set to ”connected” since it continuously streams data to mobile devices.
@@ 7.1. Impact of Sensing Different Body Parts on Activity Recognition Performance
7.2. Impact of Different Placements of the ActSonic Sensing System on the Eyeglasses Form Factor
Figure 11. Impact of active acoustic sensing placement on recognition performance of ActSonic: Left indicates a speaker (emitting C-FMCW chirps within the frequency range of 18.0-21.0 KHz) and microphone pair on the left hinge of the eyeglasses, Right indicates a speaker (emitting C-FMCW chirps within the frequency range of 21.5-24.5 KHz) and microphone pair on the right hinge of the eyeglasses, and Both indicates both aforementioned speaker-mic pairs are placed on both hinges of the eyeglasses.
We evaluated the impact of sensing system placement on the performance of ActSonic in this experiment. The raw audio data was collected in user studies using the ActSonic sensing system placed on both hinges of the eyeglasses. To measure the impact of sensor placement, we utilized the acoustic data from one side, either left or right. We then computed the echo profile and acoustic flow using that data. With only one side activated, we have just one C-FMCW frequency range, yielding echo profiles with one channel containing information on the direct acoustic transmission path (details of transmission paths are mentioned in Sec. 3.2). We customized the deep learning pipeline to take input sliding windows of the same shape with one channel and evaluated the activity recognition performance of ActSonic.
Figure 11 illustrates that placing the sensing system on both hinges of the eyeglasses yields the best performance, with a mean macro F1 score of $0.866\pm 0.044$ in our leave-one-participant-out evaluation. The right-side placement demonstrated moderate performance, with a mean macro F1 score of $0.838\pm 0.057$, which is close to the performance of both-side placements. This can be explained by the involvement of the dominant hand in most of the activities in the ActSonic recognition set, which was the right hand for all the participants in the study. On the other hand, the left hinge placement demonstrated degraded performance compared to the right one, with a mean macro F1 score of $0.811\pm 0.062$.
@@ 8.1. Visual Explanation of Learned Parameters of ActSonic Model through Saliency Analysis
To address this, we conducted saliency analysis using Gradient-weighted Class Activation Mapping (Grad-CAM) (Selvaraju et al., 2019) to visualize the ResNet18 encoder’s final convolutional layer. Overlaying feature heatmaps on acoustic flow or differential echo profile channels confirms the model’s ability to capture the movements of body parts and areas related to each activity. Figure 12 demonstrates Grad-CAM (Selvaraju et al., 2019) for all activities in the dataset.
@@ 8.3. Model Quantization and MCU Inference
To explore the feasibility of deploying our framework on glasses, we implemented the model pipeline on the MAX78002 microcontroller unit (MCU), leveraging its built-in ultra-low-power CNN accelerator. Initially, we quantized the model by converting high-precision floating-point model parameters to 8-bit integers, a necessary step for deployment on the MAX78002 MCU. Subsequently, we generated a C program for the quantized model inference using the ai8x (AI, [n. d.]) library provided by the MCU manufacturer. Due to hardware constraints, certain adjustments were made to the model pipeline for compatibility. Notably, 2D convolution kernel sizes were limited to $(1\times 1)$ or $(3\times 3)$, with fixed stride size at $(1\times 1)$. Additionally, the fully connected layer was capped at a maximum of $1024$ input neurons on the chip.
Although we successfully ran the ActSonic model on the MAX78002 MCU, the computation of the echo profile and inference of one sliding window through the quantized ResNet18 model on the MCU resulted in a mean latency of 1085.54 milliseconds or 1.09 seconds, yielding a refresh rate of 0.92 Hz. Since we predict activity labels for each 2-second sliding window with 50% overlap, the resulting refresh rate is slower than what is required for real-time inference. The major bottleneck in this process is the computation of the echo profile from the received signal, as this process is not a neural network operation supported by the MAX78002 MCU accelerator. This issue can be alleviated by transforming the cross-correlation operation into an MCU-compatible and accelerated convolution operation, which we leave for future work.
@@ 8.4. Robustness to Ambient Noise
The ActSonic system utilizes active acoustic sensing to detect daily activities by monitoring body motions. We evaluated its resilience to environmental noise across $19$ participants’ homes during the studies. Despite varying environmental factors such as HVAC, running water, TV, and ambient sounds (as detailed in Sec. 5), ActSonic maintains consistent performance independent of environmental settings. This resilience stems from its reliance on ultrasonic frequencies ($18$ kHz to $24.5$ kHz), which exceed most environmental noise sources (recorded at frequencies below $7.5$ kHz).
To confirm this hypothesis, we conducted a noise injection experiment. In this experiment, we overlaid noise from five sources commonly found in households onto the acoustic data recorded with ActSonic’s sensing module. These noise sources are people talking ($61.1$ dB), TV/radio ($68.8$ dB), music ($71.5$ dB), pets (dogs barking at $82.1$ dB and cats meowing at $51.0$ dB), and kitchen appliances (microwave at $60.5$ dB and blender at $88.7$ dB). We then computed the echo profile (as described in Sec. 3.2) and fed it into the ActSonic deep learning framework to evaluate the impact of environmental noise.
To evaluate the impact of different noises on the results, we conducted a one-way repeated measures ANOVA. The $F$-ratio was computed as $0.526$, with degrees of freedom between groups being $5$ and degrees of freedom within groups being $108$. The $p$-value was measured as $0.13>0.05$, indicating no statistically significant difference between raw and noise-injected acoustic data for tracking everyday activities using the ActSonic system. This lack of significant difference can be attributed to the bandpass filter in the ActSonic pipeline, which cancels out the audible frequency range where the injected noises occur. Nevertheless, further investigation into the impact of ambient noise is crucial to evaluate performance under real-world usage scenarios, which we leave for future work.
8.5. Effectiveness under Multi-user Scenario
Although ActSonic is a wearable system designed to track the everyday activities of the person wearing the eyeglasses, the system can be influenced by the movements of other people and objects surrounding the user, as it relies on the reflected acoustic signals of the transmitted wave. The input to the ActSonic deep learning pipeline is restricted to a sensing range of 101.185 cm, and theoretically, it should not contain any information from reflected signals beyond this range.
In the 19-person user study conducted at participants’ homes to evaluate the efficacy of the ActSonic activity recognition framework, we did not impose any restrictions on the number of other people who could be present at home during the study. To further evaluate the multi-user scenario, we conducted a study with 3 participants from the research team. In this evaluation, each person wore the device and performed 21 activities from the ActSonic tracking set (excluding bathroom activities, detailed in Sec. 5.1). While one person performed the activities, the other two were present in the same room, both moving around and staying in place randomly. Additionally, a pedestal fan with a diameter of 16 inches and a height of 47 inches was randomly turned on and off in the same room.
Our leave-one-participant-out evaluation on these 21 activities yielded a mean macro F1 score of $0.857\pm 0.029$ across the three users, which is close to the accuracy of the ActSonic system evaluated in the 19-person user study. However, we did not evaluate the scenario where multiple users in the same room each wore a ActSonic prototype. We leave this evaluation for future investigation.
8.6. Health Implication and Usability
ActSonic emits FMCW-encoded ultrasonic waves for active acoustic sensing. To assess health implications, we measured the transmitted signal intensity using a CDC-provided mobile app (Chucri A. Kardous and Peter B. Shaw, [n. d.]). The resultant intensity is $68$ dB(A), well below the $85$ dB limit set by NIOSH (Murphy and Franks, 2002). Research (Moyano et al., 2022) on MHz-range ultrasonic exposure suggests muscle tissue discomfort. However, ActSonic operates in the KHz range just above the audible threshold, with no reported issues in this range.
A major concern of active acoustic sensing systems is frequency leakage into the audible range, potentially creating noise from the transmitter. To evaluate this, we collected feedback from participants through a questionnaire after the study. All 19 participants reported that they did not hear any noise or audible sound from the sensing system while wearing the ActSonic form factor. Additionally, none reported any comfort issues. However, two participants mentioned that the glass frame size did not fit their faces, four reported needing their prescription lenses while performing the activities, and one reported that the chest-mounted camera interrupted activities such as eating and reading.
Future investigation should focus on building an attachable sensing module that can be used with any glass frame and integrating the ground truth collection module as an egocentric camera into the eyeglasses form factor. Future studies will also explore potential audibility among animals and children despite its inaudibility to adults for long-term usage.
@@ 8.8. Potential Real-world Application
8.8.1. High-resolution Behavior Data in the Wild for Health Monitoring
ActSonic can be used to journal various everyday activities for different purposes. For instance, journaling food intake behavior is crucial in combating eating disorders, often recorded manually. Previous systems required multiple sensors, had low time resolution (e.g., recognizing a meal every 10 minutes) (Shin et al., 2022), or needed training data from a user. In contrast, ActSonic can recognize eating moments at 1 Hz with over 90% F1 score in real-world settings without needing training data from a new user, facilitating immediate deployment for eating journaling practices.
Additionally, ActSonic can track 27 everyday activities, many related to health behaviors. Automatically logging these activities can help researchers and clinicians better understand a user’s activities in the wild. For instance, eating and drinking behavior can be analyzed for eating disorders and hydration levels, while activities like brushing teeth, flossing, and rinsing the mouth can be tracked with low error rates, aiding dentists in monitoring patient dental behavior.
@@ 9. CONCLUSION
This paper introduces ActSonic, a low-power and unobtrusive action recognition system that employs acoustic sensing on smart glasses. Extensive experiments involving $19$ participants in real-world settings showcase ActSonic’s adeptness in distinguishing a diverse range of everyday actions across different environments, achieving a mean accuracy of 86.6% in unconstrained user-independent evaluations. We envision ActSonic as a straightforward and efficient supplementary modality for egocentric action recognition, addressing concerns regarding privacy.
ACKNOWLEDGMENTS
This project was supported by the National Science Foundation Grant No. 2239569 and partially by the Cornell University IGNITE Innovation Acceleration Program. ChatGPT was used to polish the writing of the paper.